OpenStudy (anonymous):

given the function f(x)=x^2-x+7 find each of the following f(1)= f(-10)= f(0) =

OpenStudy (saifoo.khan):

insert the values instead of x for all three.

OpenStudy (anonymous):

how?

OpenStudy (saifoo.khan):

i will do the first one.

OpenStudy (anonymous):

okay is the first one 8?

OpenStudy (saifoo.khan):

f(x)=x^2-x+7 f(-1)=(-1)^2-(-1)+7

OpenStudy (saifoo.khan):

1+1+7 9

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the second part is 17?

OpenStudy (anonymous):

or is it 3?

OpenStudy (saifoo.khan):

117

OpenStudy (saifoo.khan):

100+10+7

OpenStudy (anonymous):

The second part is 117?

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

so its 9,3 , and 7 ...correct?

OpenStudy (anonymous):

?

OpenStudy (saifoo.khan):

9, 3, 7??

OpenStudy (anonymous):

That s the answers for f(1)= f(-10)= f(0)=

OpenStudy (anonymous):

u there .. i m taking a test and my time is running out